OpenStudy (anonymous):

looking for help with continuously compounded interest... how do you find initial investment and annual percent rate given the time to double is 7 yrs and amount after 10 yrs is 2400?

OpenStudy (adamaero):

Some more good info: http://www.harpercollege.edu/mhealy/ecogif/uein/rule70.gif

OpenStudy (paxpolaris):

\[\Large A=P e ^{rt}\]

OpenStudy (anonymous):

i know the formula but I'm not sure how to find it

OpenStudy (paxpolaris):

doubles in 7 years :\[\Large 2P = Pe^{r \cdot 7}\] solve for r

OpenStudy (anonymous):

so r=9.9%?

OpenStudy (paxpolaris):

right

OpenStudy (anonymous):

thx so much!

OpenStudy (anonymous):

how do i find the initial?

OpenStudy (paxpolaris):

amount after 10 years is 2400 just plug in: A=2400 t=10 r= 0.09902 and solve for P

OpenStudy (anonymous):

i divided 2400 by e^9.9 and it gave me a decimal

OpenStudy (paxpolaris):

e^0.99

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

Determine the time necessary for P dollars to double when it is invested at interest rate r compounded monthly, daily. r =3.4%

OpenStudy (anonymous):

sorry I'm so lost;/

OpenStudy (adamaero):

except you have to do extra since it's not a year but \[\frac{ 1 }{ 12 }th\] of a year and \[\frac{ 1 }{ 52 }\] of a year.